During a test flight, an aircraft flies from Sandy Land to White City and back to Sandy Land. The distance between Sandy Land and White City is 450 km and the total time taken for the whole journey is 5 hours and 30 minutes. Given that there is a constant wind blowing from Sandy Land to White City and that the speed of the aircraft in still air is 165km/h, find the speed of the wind. State the assumptions you have made to this problem.
Hint : Let the speed of the wind be x km/h.

Answer by jorel1380(3719)

The distance is 450 km between the two cities. If the speed of the wind is x kph, then, assuming the wind is constant:
450/165+x + 450/165-x=5.5 (5-1/2 hours)
450(165-x)+450(165+x)=5.5(27225-x²)
Solve for x
☺☺☺☺